OpenStudy (anonymous):

3^x - 2^(x-1) = 3^(x-2) + 2^(x+1) <=> 3^x - 3^(x-2) = 2^(x+1) + 2^(x-1) <=> (3^x)(1-1/9) = (2^x)(2+1/2) <=> (3/2)^x = (8/9) / (5/2) <=> (3/2)^x = 45/16 Am I mistaken? Is there anything else to calculate the solution?

OpenStudy (amistre64):

i assume your solving for x?

OpenStudy (anonymous):

Yes!

OpenStudy (amistre64):

3^x - 2^(x-1) = 3^(x-2) + 2^(x+1) 3^(x) - 2^(x) a = 3^(x) b + 2^(x) c 3^(x) - 3^(x) b = 2^(x) c + 2^(x) a 3^(x) (1-b) = 2^(x) (c+a) logs help when working with exponents so , log[3^(x) (1-b)] = log[2^(x) (c+a)] log[3^(x)] + log[(1-b)] = log[2^(x)] + log[(c+a)] x log[3] + log[(1-b)] = x log[2] + log[(c+a)] x log[3] - x log[2] = log[(c+a)] - log[(1-b)] x ( log[3] - log[2]) = log[(c+a)] - log[(1-b)] thats my thoghts

OpenStudy (anonymous):

Yep, I see! I am supposed to do so without logarithms though as I will learn about them in the next chapter..

OpenStudy (amistre64):

without logs? then some deformed rt :)

OpenStudy (amistre64):

maybe ... 3^(x) (1-b) = 2^(x) (c+a) (3/2)^x = (c+a)/(1-b) trial and error if the left side plays nice

OpenStudy (amistre64):

a=1/2, b = 1/9, c = 2 3*9 3*3*3 --- = ------- 2*8 2*2*2*2 almost .... as is i dont see a 'simple' solution since the number of 3s does not equal the number of 2s

OpenStudy (anonymous):

Thats what I ended up with too so I guess thats it..

OpenStudy (amistre64):

thats the best i can do :) good luck